OpenStudy (anonymous):

Q. a swimmer goes from A to B in 4 hrs and returns back from B to A in only 1 hr.find velocity of river flow and velocity of swimmer in still water.

OpenStudy (anonymous):

x=v*t x=(v(swimmer)-v(river))*4 x=(v(swimmer)+v(river))*1 (v(swimmer)+v(river))*1=(v(swimmer)-v(river))*4 v(s)+v(r) = 4v(s) - 4v(r) 5v(r) = 3v(s) to find the velocity, you should know what is distance.

OpenStudy (matthewrlee):

That's a dumb swimmer. Who swims from A to B when it takes 4 hours? That's why we have boats. He got back to A three hours early cause someone pulled him out of the water and stuck him in a boat. That's want happened.

OpenStudy (anonymous):

well the first answer is almost there..to put in other words..his results mean that for the information given it could be an infinite number of solutions..as long as the the rivers speed is 3/5 of the speed of the swimmer ( so a river speed of 3 mph and a swimmers speed of 5 mph would work for example). Unless they tell you the distance traveled or the speed of the river you will not be able to solve for one specific value of the swimmer s speed.

OpenStudy (anonymous):

To solve this problem, you can set the distance AB to an arbitrary value since the swimmer is traveling the same distance each time. In this case I set it to 1 km since that is easiest. From there a substitution of equations will give you the two resulting velocities, as long as you are careful with your signs. Please see the attachment for the full solution.